OpenStudy (anonymous):

Find the exact value of (-1/2 + i(sqrt(3)/2))^7 and write the result in exact rectangular form.

OpenStudy (anonymous):

\[(-\frac{ 1 }{ 2 } + i \frac{ \sqrt(3) }{ 2 })^{7}\]

OpenStudy (anonymous):

@RadEn, do you know?

OpenStudy (raden):

actually, use the binomial newton formula : (a+b)^n = nC0(a)^(n-0)(b)^0 + nC1(a)^(n-1)(b) + nC2(a)^(n-2)(b)^2 + ... + nCn(a)^(n-n)(b)^n +

OpenStudy (anonymous):

I have not learned that

OpenStudy (raden):

hmmmm... for basic, u have to know about factorial like this : 4! = 4*3*2*1 = 24 3! = 3*2*1 = 6 and to calculating for combination, u can use the formula : nCr = n!/((n-r)!r!)

OpenStudy (raden):

maybe this site can help u http://www.mathsisfun.com/pascals-triangle.html

OpenStudy (anonymous):

I couldn't figure it out. This was on my final and I just took it. It was one of the extra credit problems so i'm not worried about it. Thanks anyway!